OpenStudy (anonymous):

Differentiate the function y = f(x) = ln(tan(x)).......... is the correct answer cot(x)sec^2x by chain rule? Differentiate the function y = f(x) = ln(tan(x)).......... is the correct answer cot(x)sec^2x by chain rule? @Mathematics

OpenStudy (anonymous):

f(x) = ln(tan(x)) df/dx = 1 / sin *1/cos = sec * csc